The police notice 5 drunken drivers daily in average with the variance 4, independently everyday. What is the probability that they catch over 1900 drunken drivers yearly?

Answer by rothauserc(4718)

1900/365 = 5.2 daily
:
assume a normal distribution
:
z-score = (5.2 - 5) / 4 = 0.05
:
checking table of z-scores, we find
:
Probability (P) ( X < 5.2 ) = 0.5199
:
P ( X > 5.2 ) = 1 - 0.5199 = 0.4801 approximately 0.48
: